# NeuroBench: A Framework for Benchmarking Neuromorphic Computing Algorithms and Systems

### v3

@@ Results
    NeuroBench will provide a QUBO generator that can uniquely specify each workload by three specific parameters provided by the benchmark: the number of graph nodes, the sparsity of graph connections, and a random seed. The generator provides a large dataset for reliable statistics and allows scaling from modest workloads for small-scale and prototype systems to large workloads for larger-scale systems. An independent dataset is provided to tune the hyperparameters of the solver. The SUT is evaluated based on three metrics: The first success metric is the maximum supported workload size. The second and third metric is the time and energy required to obtain pre-defined levels of optimality, respectively. The solution optimality is measured by the size of the independent set of nodes found by the SUT.
System Track Harness
Figure 7: An overview of the NeuroBench harness supporting the system track.
A diagram of the NeuroBench harness with extended support for the system track is shown in Figure 7. Blue boxes in the figure denote hardware-specific infrastructure components which will connect to the general top-level harness interfaces. Common interfaces within the runtime enforce that the harness can be modularly extended to various system backends.
Within the system runtime, intermediate representations (IR) are used to compile and map models. In ML development, the ONNX IR [63] has been an enabler of cross-platform portability, thus IR infrastructure within the harness is an important component of common tooling. A common high-level IR such as the Neuromorphic Intermediate Representation (NIR) [64] or Lava [36] can unify spatio-temporal graphs describing the neuromorphic models, and may be transformed into a more optimized low-level IR that is specific to the target hardware (e.g., sPyNNaker machine graph [65]). The inclusion of shared IRs within the harness allows for benchmarking equivalent models across multiple systems, which will highlight system optimizations for cross-compatibility. The IRs may also be hardware-specific, tightly integrated within the system to maximize performance.
From a top-level API, users will be able to run algorithm benchmarks or system benchmarks through the corresponding runtimes. By providing the runtimes side-by-side, the harness presents a single tool for both algorithm and system benchmarking. As the algorithm and system tracks mature to support the same tasks, the harness will accelerate benchmarking by facilitating quick prototype complexity analysis followed by deployed performance measurement with minimal implementation overhead.
@@ Discussion
Benchmarking neuromorphic computing has faced challenges stemming from the diversity of neuromorphic approaches, the range of implementation and deployment tools, and rapid research evolution. NeuroBench addresses these challenges as a framework for the inclusive, actionable, and iterative benchmarking of neuromorphic solutions, by including novel tasks and metrics, open-source and extendable harness tools, and facilitating systematic growth via community collaboration. NeuroBench is supported and developed by a broad community of neuromorphic researchers to be a standard, agreed-upon benchmarking framework for neuromorphic technology.
Initial NeuroBench benchmarks span applications across domains of continual learning, computer vision, sensorimotor prediction, and time-series forecasting, as well as system implementations for audio and optimization settings. Baselines for each benchmark of the complete v1.0 algorithm track demonstrate the utility and validity of the metric framework, and offer a starting point to further algorithmic research in model architecture and training for greater performance and lowered complexity.
The initial NeuroBench algorithm and system tracks achieve the first steps of designing benchmarks for algorithms executed in a digital time-stepped fashion and systems in mature deployment stages. In order to expand inclusive and fair benchmarking to further approaches, the next milestones of the NeuroBench project will be to extend metrics and standard protocol designs to cover continuous-time execution and a wider range of hardware platforms including FPGAs, custom integrated circuits, as well as more exploratory platforms in simulation stage, such as memristive hardware.
@@ Methods
Specifications of the Algorithm Track Metrics
@@
Execution rate is a metric which is not directly computed by the harness, but should be reported by the user. The metric reflects the real-time correlation of the rate at which the model computes input data. If the model processes input with a temporal stride of $t$ seconds, then the rate should be reported as $t^{-1}$ Hz. Note the distinction between stride and bin window - input can be binned in overlapping windows, but execution rate depends on the temporal stride of window processing. As an example, a model may use 50 ms windows of input and compute every 10 ms, which would give an execution rate of 100 Hz.
This metric is currently not well-defined for models operating under event-based or continuous-time contexts. These limitations will be addressed in future benchmark versions.
@@
For each language, the longest length words (that had the appropriate number of samples) were selected to allow for rich and robust temporal features to be learned. Next to the richness of longer words, there are practical considerations for this choice. The MSWC dataset normalizes all samples to a duration of 1 second, centered around the 0.5 seconds mark. For shorter words, this means that the data needs to be zero-padded on both sides to fill the entire duration. Longest-length words are likely to fill the complete sample and reduce zero-padding, which is also useful in scenarios in which algorithms seek to classify words before the sample has completed [67]. Furthermore, common keyword spotting solutions, such as Ok Google, Alexa, and Hey Siri, use multi-syllable wake-phrases to assist in accurate word classification.

### v5

@@ Introduction
    NeuroBench provides a QUBO generator that can uniquely specify each workload by three specific parameters provided by the benchmark: the number of graph nodes, the density of graph connections, and a random seed. The generator provides a large dataset for reliable statistics and allows scaling from modest workloads for small-scale and prototype systems to large workloads for larger-scale systems. The graph sizes specified by the benchmark increase in a pseudo-geometric progression (10, 25, 50, 100, 250, …), and submissions are encouraged to extend the problem size to the limits of the SUT. Graph density ranges from 1% to 30%, in order to show the relationship between the number of connections and SUT power. 5 random seeds for each setting should be tested.
    Optimization algorithms use heuristic methods to iteratively refine approximate solutions to intractable problems that cannot be completely solved. The QUBO benchmark thus measures solution optimality and energy consumption after fixed, pre-set runtimes, removing any timing measurement. Solution optimality is defined as BKS-Gap, a relative gap between the current SUT’s solution compared against the best-known solution (BKS) to the same problem found using a high-powered solver with a long runtime.
Baseline Results
Baseline results for each of the two system track benchmarks are provided for a mature neuromorphic system against a conventional platform. Like for the algorithm track baseline results, the system track baselines are intended to snapshot the solution space and provide starting points for the task leaderboards. Further details on each baseline system are available in the Methods section, and in-depth system documentation for the Xylo ASC baseline and CPU/Loihi 2 QUBO baselines are provided by Ke et al. [72] and Pierro et al. [57], respectively.
Acoustic Scene Classification
For the acoustic scene classification task, two baseline embedded systems are reported:
  • CPU – The CPU baseline is an Arduino Nano 33 BLE, which uses an ARM Cortex M4 microcontroller for both pre-processing and inference. The digital audio sample is pre-processed using Mel-filterbank energies (MFE), and inference uses a conventional CNN. Execution time is measured using on-chip timers, and power is measured using total system power.
  • Xylo – The Synsense Xylo [73] neuromorphic baseline uses a feed-forward SNN with multiple synaptic time constants [54]. As the system is intended for continuous, real-time audio processing, the board uses an analog front end to pre-process analog audio signals directly from a microphone into spikes for the digital inference engine. To conform with a digital benchmark dataset, a simulator of the analog pre-processor generates spikes, which are routed to the inference module. Execution time and power are measured using on-board instruments.
Baseline Accuracy Execution Time (ms) Idle Power (mW) Active Power (mW) Dynamic Power (mW) Dynamic Energy (mJ/inf) CPU pre-process 79.64% 43 79.40 100.72 21.32 0.917 (system-wise) inference 45 79.40 100.15 20.75 0.934 Xylo pre-process* 79.90% - 0.00017* 0.015* 0.015* 0.015* (component-wise) inference 84 0.351 0.692 0.341 0.028
Table 6: Baseline results for the acoustic scene classification task. Pre-processing of the Xylo (marked with an asterisk *) measures power of the analog pre-processor in real-time relative to the audio data, whereas other measurements are of digital components processing digital data on-hand. Idle Xylo pre-processing measures silence and active measures test data audio played to the device, and energy is measured over the sample duration of 1 second. CPU power is measured over the full Arduino system, as it does not have on-board power instrumentation, while Xylo power measures power consumed by the Xylo Audio 2 ASIC only. Dynamic power and energy provide proper comparison between the systems, and idle and active measurements are provided for transparency.
Table 6 lists baseline results for the neuromorphic Xylo system against an Arduino system. Compared to prior neuromorphic audio system benchmarking [39], which takes a server-class CPU as a point of comparison, we adopt a fairer approach by focusing on low-power edge application and comparing against an Arduino embedded microprocessor. At comparable inference accuracy, Xylo exhibits 60.9×60.9$\times$60.9 × less dynamic inference power and 33.4×33.4$\times$33.4 × less dynamic inference energy consumption than the Arduino.
QUBO
Figure 7: Percentage gap from the best known solution (BKS-Gap%) for the QUBO workloads with QUBO matrices at $15\%$ density (lower is better). Results are shown for different timeouts of the QUBO solvers. Figure taken, with permission, from Pierro et al. [57].
Figure 8: Power consumption of the QUBO solvers running simulated annealing (SA) or TABU search on CPU, and the parallelized version of simulated annealing on Loihi 2. The CPU solvers require up to 37×37$\times$37 × more power than the neuromorphic algorithm on Loihi 2. Since the QUBO workloads were run for a fixed timeout, differences in power consumption are equivalent to differences in energy consumption of the processors. Figure taken, with permission, from Pierro et al. [57].
Three baselines are measured for the QUBO benchmark:
  • Simulated Annealing (SA) – The simulated annealing solver uses Markov chain Monte Carlo (MCMC) sampling to probabilistically explore the search space.
  • Tabu Search (TABU) – Tabu search solvers maintain and iterate a list of prohibited actions in order to prevent the search from remaining in local minima or revisiting states. Both the TABU and SA solver baselines use the D-Wave Samplers library [74] on an Intel Core i9-7920X desktop-class CPU, with power measured using Intel SoC Watch.
  • Loihi 2 – The Loihi 2 neuromorphic system solver uses an SNN formulation of the simulated annealing algorithm which enables solving via neural dynamics, and parallelization via stochastic refractory periods. The baseline is implemented on one Loihi 2 chip on the 8-chip Kapoho Point board, and internal power instrumentation measures all compute and memory components of the chip.
Figure 7 shows the optimality reached by the CPU- and Loihi 2-based solvers after different timeouts. For tight time constraints, at $10^{-2}$ seconds timeout or less, Loihi 2 finds feasible solutions to workloads 4×4$\times$4 × larger than the CPU. But for timeout lasting $10$ seconds or longer, the CPU running TABU provides the lowest BKS-Gap, incentivizing algorithmic advances for neuromorphic optimization systems. Figure 8 illustrates the power consumed during runtime. Across the workloads, the Loihi 2 solver requires 37.24×37.24$\times$37.24 × less power compared to the best CPU solver.
Discussion and Future Work
The initial baselines for the v1.0 system track compare correctness, timing, and efficiency of neuromorphic systems against conventional CPU systems in domains of both audio classification and optimization. Against mature, commercially-developed CPU systems, for both edge and server use cases, the neuromorphic systems show strong advantages in general efficiency, as well as further promises in terms of timing and correctness.
In future NeuroBench iterations, the system track benchmarks can be unified under common tooling, similar to the algorithm track. Software toolchains such as Lava [36], Fugu [37], SPyNNaker [75], and Samna [76], among others, have been developed to interface with specific hardware platforms. Many of the stacks are built with general paradigms to support extension to any backend, and the community is actively moving towards developing standards for deployment tools. The current v1.0 benchmark specifications allow for open algorithm and software design in order to demonstrate fully optimized performance for neuromorphic systems. As standards mature in the future, a core focus of the NeuroBench system track is to introduce a closed-algorithm benchmarking category that leverages the recently proposed NIR model description framework [77] as a general, cross-platform tool for benchmarking key workloads of interest across many different platforms.
@@ Discussion
Benchmarking neuromorphic computing has faced challenges stemming from the diversity of neuromorphic approaches, the range of implementation and deployment tools, and rapid research evolution. NeuroBench addresses these challenges as a framework for the inclusive, actionable, and iterative benchmarking of neuromorphic solutions, by including novel tasks and metrics, open-source and extendable harness tooling, and facilitating systematic growth via community collaboration. NeuroBench is supported and developed by a broad community of neuromorphic researchers to be a standard, agreed-upon benchmarking framework for neuromorphic technology.
Initial NeuroBench benchmarks span applications across domains of continual learning, computer vision, sensorimotor prediction, and time-series forecasting, as well as system implementations for audio and optimization settings. Baselines for each benchmark of the complete v1.0 algorithm track demonstrate the utility and validity of the metric framework, and offer a starting point to further algorithmic research in model architecture and training for greater performance and lowered complexity. The system track v1.0 benchmark baselines similarly provide the groundwork for both conventional and head-to-head comparisons of deployed hardware systems. Via collaboratively-defined measurement methodology and task specifications, the shared system track guidelines will be modified and extended to define standard protocols for continuous-time execution, analog circuit implementations, and other exploratory platforms in simulation stages, such as memristive hardware.
@@ Methods
Algorithm Track Metrics
@@
Execution rate is a numeric which is not directly computed by the harness, but should be reported by the user. The numeric reflects the real-time correlation of the rate at which the model computes input data. If the model processes input with a temporal stride of $t$ seconds, then the rate should be reported as $t^{-1}$ Hz. Note the distinction between stride and bin window - input can be binned in overlapping windows, but execution rate depends on the temporal stride of window processing. As an example, a model may use 50 ms windows of input and compute every 10 ms, which would give an execution rate of 100 Hz.
This numeric is currently not well-defined for models operating under event-based or continuous-time contexts. These limitations will be addressed in future benchmark versions.
@@
For each language, the longest length words (that had the appropriate number of samples) were selected to allow for rich and robust temporal features to be learned. Next to the richness of longer words, there are practical considerations for this choice. The MSWC dataset normalizes all samples to a duration of 1 second, centered around the 0.5 seconds mark. For shorter words, this means that the data needs to be zero-padded on both sides to fill the entire duration. Longest-length words are likely to fill the complete sample and reduce zero-padding, which is also useful in scenarios in which algorithms seek to classify words before the sample has completed [79]. Furthermore, common keyword spotting solutions, such as Ok Google, Alexa, and Hey Siri, use multi-syllable wake-phrases to assist in accurate word classification. Using shorter keyword subsets can lead to greater challenge in both base language training and continual class learning. The full list of chosen words for the task presented in this paper (longest words), as well as other potential subsets of short words for the same languages, can be found with dataset documentation through the harness.
@@
System Track Metrics
Given the variability of task application areas and system sizes, the NeuroBench metric methodology is defined on a per-benchmark basis. Particularly for efficiency metrics, as neuromorphic systems are not matured, a singular power measurement method cannot be applied to all submissions. It is the responsibility of the submitter to faithfully capture all active processing components for their system and fairly outline their methodology in the report. As NeuroBench moves towards head-to-head benchmarking of neuromorphic systems by hardware vendors and owners, official results must be associated with a report that provides context for the benchmark submission results, as the overall benchmark format is generally open and does not have stringent consistency rules. The report must contain
  • an outline of the system architecture,
  • an outline of the algorithm used (model architecture, tuning),
  • a diagram depicting the workflow, including (where applicable) data initialization, host pre-processing, data loading, on-device preprocessing, inference, post-processing,
  • timing measurement description,
  • power measurement description, including measurement devices, included hardware components, and measurement time resolution,
  • a re-iteration of the results.
In addition, official submissions will be subject to potential audits during which an auditor will inspect the methodology and request additions or revisions to the results and report if necessary.
System Track Benchmarks
Detailed information on each v1.0 system track benchmark is provided here, and the most updated information can be found in the official NeuroBench system track documentation: https://github.com/NeuroBench/system_benchmarks
Acoustic Scene Classification (ASC)
Benchmark Dataset
The dataset is based on the DCASE 2020 acoustic scene classification challenge [68], using the TAU Urban Acoustic Scenes 2020 Mobile datasets. Of the 10 available scene classes, 4 are used: “airport”, “street_traffic”, “bus”, “park”. Of the 9 real / simulated audio recoding devices available, 1 real device is used: “a”. Audio samples are sliced into 1-second samples. The audio may be resampled to a different frequency as a pre-processing step, which is not included in inference measurement. 41360 training samples are available, as well as 16240 test samples. The NeuroBench system track repository linked above provides a download script and a PyTorch-compatible dataset file which is expected to be used as the front-end data generator for all submissions. The data may be reformatted into a different framework, this is not included in inference measurement.
Task
After training a model using the training set, the submitter will report test set accuracy, as well as execution time and energy per inference on the system. The test split should not be used to train or tune the model, only the train split. Audio samples will be processed in batch-size 1, in which one sample is processed at a time and the next sample is not processed until the previous one is finished. The general compute flow consists of three steps: pre-processing, inference, and classification. One inference is defined as the processing of one second of audio data. Inference is separate from classification because systems are often intended to process samples as sequences over time, rather than all at once, and the classification may be available before the whole data sequence is seen. Classification thus does not need to be included in the benchmark measurement. Pre-processing in general can be defined as feature extraction or the conversion of raw audio data into a format that is suitable for inference. Execution time and energy of pre-processing must be included in benchmark measurement, and will be reported separately from inference. In certain systems, e.g. Synsense Xylo, pre-processing blocks use analog hardware on-chip to directly convert real-time analog microphone output to digital spikes for inference. In order to operate inference from a digitally-encoded dataset, this pre-processing must be simulated and the spikes are sent through a side-channel directly to the inference block. For such cases, the reported pre-processing measurements will be for the analog hardware running in real-time on the dataset audio played over speakers into the device microphone.
Metrics
The following metrics should be reported:
  • Accuracy – Accuracy of the predictions on the test set, measured from the system and not in any software simulation.
  • Execution time – Execution time is the average time per sample for pre-processing and inference. The final result should be averaged over all samples in the test set. The time begins when data has been loaded into on-board or on-chip memory and is prepared to be processed. The time ends when the last timestep of the sequence has completed processing.
  • Power/Energy – Idle power and active power should be reported, where idle power is power of the chip when it has been configured and it has prepared to begin processing. Note that this should not measure the device in a lower-power sleeping state. Active power is the average power consumed during processing. The difference between active power and idle power should be reported as dynamic power. Power should be converted to energy by multiplying power by the averaged execution time. The power measurement should include all computational components and power domains which are active during the workload processing. If applicable, power may be measured over longer data sequences (e.g. 5 seconds rather than 1 second), such that configuration costs are amortized over a longer period of processing. This should be done separately from accuracy and execution time experiments, and should be clearly detailed in the associated submission report.
Quadratic Unconstrained Binary Optimization (QUBO)
Quadratic Unconstrained Binary Optimization (QUBO) refers to the problem of finding the binary variable assignment $x_{i}\in\{0,1\}$ that optimizes the quadratic cost function
$\min_{\mathbf{x}\in\{0,1\}^{n}}c(\mathbf{x})=\min_{\mathbf{x}\in\{0,1\}^{n}}\mathbf{x}^{T}\mathbf{Q}\mathbf{x}$
subject to no constraints.
The solvers for QUBO will be benchmarked using Maximum Independent Set (MIS) workloads. Given an undirected graph $\mathcal{G}=(\mathcal{V},\mathcal{E})$, an independent set $\mathcal{I}$ is a subset of $\mathcal{V}$ such that, for any two vertices $u,v\in\mathcal{I}$, there is no edge connecting them, i.e., $\nexists\;e\in\mathcal{E}\;s.t.\;e=(u,v)\;\vee\;e=(v,u)$.
The MIS problem has a natural QUBO formulation: for each node $u\in\mathcal{V}$ in the graph, a binary variable $x_{u}$ is introduced to model the inclusion or not of $u$ in the candidate solution. Summing the quadratic terms $x_{u}^{2}$ will thus result in the size of the set of selected nodes. To penalize the selection of nodes that are not mutually independent, a penalization term is associated to the interactions $x_{u}x_{v}$ if $u$ and $v$ are connected. The resulting $\mathbf{Q}$ matrix coefficients are defined as
$q_{uv}=\begin{cases}-1&\text{if }u=v\\ 4&\text{if }u\neq v\text{ and }(u,v)\in\mathcal{E}\\ 0&\text{otherwise.}\end{cases}$
The MIS problem is NP-hard and intractable, and therefore any solver system approximates a solution. Therefore, the cost of the QUBO formulation ($\mathbf{x}^{T}\mathbf{Q}\mathbf{x}$) is used to assess solutions, and solutions are not restricted to being maximum nor independent sets. The QUBO formulation ensures that any non-independent set will always have a higher cost than a corresponding independent set with conflicting nodes removed.
Benchmark Dataset
The benchmark’s workload complexity is defined such that it can automatically grow over time, as neuromorphic systems mature and are able to support larger problems.
  • Number of nodes, spaced in a pseudo-geometric progression: 10, 25, 50, 100, 250, 500, 1000, 2500, 5000, …
  • Density of edges: 1%, 5%, 10%, 25%.
  • Problem seeds: 0, 1, 2, 3, 4 are allowed for tuning. At evaluation time, for official results NeuroBench will announce five seeds for submission. Unofficial results may use seeds which are randomly generated at runtime.
Each workload will be associated with a target optimality, which is the minimum cost found using a conventional solver algorithm. Small QUBO workloads with fewer than 50 nodes will be solved to global optimality, corresponding to the true maximum independent set. Larger workloads cannot be reasonably globally solved. The DWave Tabu CPU sampler [74] will be used with 100 reads and 50 restarts, and the QUBO solution with the best cost (best-known solution, BKS) found will set the target optimality for the tuning workload seeds. For evaluation workload seeds, the same method will be used to set the target optimality. NeuroBench will provide target optimalities for workloads up to 5000 nodes. Submissions are encouraged to continue scaling up the workload size along the pattern to demonstrate the capacity of their systems. The first group that tackles workloads of an unprecedented size should provide the benchmark solutions via a pull request to the system track repository. The dataset workload generator and scripts for the DWave Tabu sampler to compute optimal costs are available in the repository, and this code is expected to be used as the front-end data generator for all submissions.
Task and Metrics
Based on the BKS found for each workload, the BKS-Gap optimality score of the solution found by the SUT is defined as
[…]
where $c_{\mathrm{target}}$ is the QUBO cost of the BKS, and $c$ is the cost found by the SUT. This may be reported as a percentage gap by multiplying by 100. If the SUT manages to beat the BKS, then the BKS-Gap will be negative.
Given each workload, the benchmark should report the BKS-Gap of the solution found by the SUT after a fixed runtime. The time begins after the graph has been loaded into the SUT. Timeouts spread across orders of magnitude ($10^{-3}$, $10^{-2}$, $10^{-1}$, $10^{0}$, $10^{1}$, $10^{2}$). As the runtime is fixed, no measured timing metric is reported, and submissions should report average power over the duration of the runtime, as it is directly proportional to energy consumed. Importantly, the QUBO solver needs a module to measure the cost of its solutions, and this module should be considered as part of the SUT, thus its computational demand and power must be included in the benchmarking results.
In the future, a different optimization task scenario may be used for the same QUBO dataset, in which the SUT must run until it reaches a small BKS-Gap, rather than stopping after a fixed-timeout. Here, the benchmark should report latency and energy required to reach BKS-Gap thresholds, e.g., 0.1, 0.05, and 0.01. This task scenario normalizes systems to solution quality, rather than SUT runtime.
System Track Baselines
Acoustic Scene Classification (ASC)
Figure 9: Left: An overview of the benchmarking system for the Xylo. A simulator of the analog encoding module is run on a PC and streamed to the SNN inference core via USB. After inference, outputs are routed back to the PC for classification. An on-board FPGA configures and records power of the Xylo components. Right: The Xylo™Audio 2 hardware development kit (HDK), used as the SUT. The red outline marks the Xylo inference module. Figures taken, with permission, from Ke et al. [72]
CPU
The CPU baseline uses an Arduino Nano 33 BLE board, which runs preprocessing and inference on an ARM Cortex-M4F microcontroller running at 64MHz. All training and deployment uses Edge Impulse, a commercially developed ML-operations platform for tinyML [84]. The 1 second digital audio samples were converted into two-dimensional frames of Mel-filterbank energies (MFE), and inference uses a network with two convolution layers with batch normalization and max pooling. The trained model was quantized then compiled to an Arduino library using Edge Impulse EON compiler, which optimizes for memory and flash usage. Execution time is measured using on-chip timers from the Arduino, while power is measured using an external multimeter for total system power. Power was measured separately for idle, active pre-processing, and active inference by taking the average power over 60 seconds. The Arduino repeatedly computes over one sample loaded in memory, with a recording frequency of 3 Hz on a Keysight 34465A digital multimeter.
Synsense Xylo
As shown in Figure 9, the Xylo baseline used a host PC to run a simulation of the analog pre-processing unit on the benchmark dataset, which was routed into the SNN core on the Xylo SUT. Accuracy is measured by performing a maximum on the output spikes of the SNN. The SNN on Xylo is a feedforward network with three hidden layers, where each layer is connected by synapses with varying time delays. Training is done using the open-source Rockpool toolchain [85]. In order to amortize measurement overheads, power and execution time are measured over continuous streams of 10 seconds of audio, where the execution time result is divided by 10 and the power result is averaged over the duration. The measurements are made by the on-board FPGA, which operates at 12.5MHz and samples power at 1280 Hz. Accuracy is still measured on the 1-second samples. Full details are available in Ke et al. [72]
Quadratic Unconstrained Binary Optimization (QUBO)
Figure 10: A Kapoho Point board with Loihi 2 chips is connected to a host PC via ethernet. Shown here is a Kapoho Point board with 8 Loihi 2 chips, while the experiments were run on a board with 1 Loihi 2 chip. The QUBO workload is loaded onto the chip, and all processing happens on Loihi 2. On Loihi 2, the neuro-cores, in yellow, are arranged in two $8\times{}8$ grids and iteratively update the variables of the QUBO workload. Embedded CPUs, shown in blue, monitor the cost of the variable assignment. Six parallel IO ports enable 3D stacking of multiple chips for larger workloads than used here. An additional IO port provides the communication to the host PC. The energy and runtime measurements cover the computation of the Loihi 2 chip.
The processors ran repeatedly with five different initial variable assignments and seeds for different timeouts between $10^{-3}$ s to $10^{3}$ s, and for different workloads sizes of up to $1000$ variables. The neuromorphic algorithm is a parallelized version of simulated annealing that was running on a Kapoho Point board with one Loihi 2 chip, as shown in Figure 10. The Loihi 2 board was controlled using Lava 0.8.0 and Lava Optimization 0.3.0. For comparison with conventional hardware, two solvers were adopted based on simulated annealing and tabu search, as implemented in the D-Wave Samplers v1.1.0 library [74]. The library was compiled on Ubuntu 20.04.6 LTS with GCC 9.4.0 and Python 3.8.10. CPU measurements were obtained on a machine with Intel Core i9-7920X CPU @ 2.90 GHz and 128GB of DDR4 RAM, using Intel SoC Watch for Linux OS 2023.2.0. All details on the solvers, benchmarking routine, and results have been provided by Pierro et al. [57].